### v1

# AnTKV: Anchor Token-Aware Sub-Bit Vector Quantization for KV Cache in Large Language Models

Abstract
Quantization has emerged as an effective and lightweight solution to reduce the memory footprint of the KV cache in Large Language Models (LLMs). Nevertheless, minimizing the performance degradation caused by ultra-low-bit KV cache quantization remains a significant challenge. We observe that quantizing the KV cache of different tokens has varying impacts on the quality of attention outputs. To systematically investigate this phenomenon, we perform forward error propagation analysis on attention and propose the Anchor Score (AnS) that quantifies the sensitivity of each token’s KV cache to quantization-induced error. Our analysis reveals significant disparities in AnS across tokens, suggesting that preserving a small subset with full precision (FP16) of high-AnS tokens can greatly mitigate accuracy loss in aggressive quantization scenarios. Based on this insight, we introduce AnTKV, a novel framework that leverages Anchor Token-aware Vector Quantization to compress the KV cache. Furthermore, to support efficient deployment, we design and develop a triton kernel that is fully compatible with FlashAttention, enabling fast online Anchor Token selection. AnTKV enables LLaMA-3-8B to handle context lengths up to 840K tokens on a single 80GB A100 GPU, while achieving up to 3.5x higher decoding throughput compared to the FP16 baseline. Our experiment results demonstrate that AnTKV matches or outperforms prior works such as KIVI, SKVQ, KVQuant, and CQ under 4-bit settings. More importantly, AnTKV achieves significantly lower perplexity under ultra-low-bit quantization on Mistral-7B, with only 6.32 at 1-bit and 8.87 at 0.375-bit, compared to the FP16 baseline of 4.73.
[…]
1 Introduction
LLMs have attracted much attention due to their strong capability in natural language processing and other artificial intelligence tasks openai2023gpt4 ; guo2025deepseek ; team2023gemini ; tang2025the ; dong2025can . With the rapid development in recent years, LLM can manage increasingly larger context lengths ranging from hundreds of thousands to millions of tokens, and its broadening scope allows LLM to tackle more complex tasks zhao2023survey ; wang2025all ; das2025security ; wang2025agenttaxo . However, these models, which are typically built on decoder-based transformer architectures, generate tokens recursively during inference, thus leading to a large accumulation of the KV cache as the context length zhu2025oraclekv ; bai2024longbench ; liu2025comprehensive and batch size increases pope2022efficientlyscalingtransformerinference . The large memory footprint of the KV cache presents a significant challenge. For long-context inference, such as LLaMA-3’s 128K context length, the KV cache can consume several gigabytes of memory, nearly matching the model’s weights. This challenge is even greater for models like Gemini, which handle contexts up to a million tokens. Moreover, LLM-RAG systems yao2025cacheblend pre-generate and store KV caches. However, the excessive size of these caches presents a significant challenge for storage.
Figure 1: Self-attention output error when quantizing the KV cache of Mistral-7B to 1-bit.
Figure 2: The perplexity of Mistral-7B on the Wikitext2 under different bit-widths.
To address the growing KV cache, various optimization strategies have been proposed liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; liu2024retrievalattention ; zhang2024cam ; liu2024kivi ; dong2025can , with quantization emerging as a prominent solution. Quantization approaches fall into two main categories: scalar quantization (SQ) and vector quantization (VQ). SQ compresses the KV cache by mapping floating-point values to fixed low-bit representations liu2024kivi ; duanmu2024skvq ; hooper2024kvquant , but the minimum 1-bit requirement limits its compression to 1/16 on FP16. In contrast, VQ maps high-dimensional vectors to a finite codebook lingle2024transformervq ; zhang2024kv , capturing intra-vector correlations. This enables higher compression rates and better preserves representational fidelity, especially under ultra-low-bit regimes.
During the LLM inference, different tokens have different effects on the accuracy of LLMs, so it is unreasonable to treat all tokens equally in VQ methods. To illustrate this, we quantize the KV cache of Mistral-7B mistral to 1-bit using VQ, and Figure 1 shows the impact of KV quantization on the self-attention output error within layer 31. It reveals that even in the absence of tokens with exceptionally large quantization errors, a few tokens (Anchor Tokens) can still introduce substantial deviations in the output of the self-attention operator. Therefore, measuring the importance of tokens during inference can further explore the performance potential of VQ methods for compressing the KV cache.
To this end, we introduce Anchor Score (AnS), a metric derived from forward error analysis, to quantify each token’s impact on the output. Leveraging AnS, we propose AnTKV, a novel framework that employs anchor token-aware VQ for KV cache compression. AnTKV mitigates quantization-induced degradation through a two-stage design: offline token-aware centroids learning and online anchor token selection. In the offline stage, the framework performs the token-aware weighted k-means clustering to generate centroids. During the online stage, it computes the AnS for each request to identify anchor tokens that significantly contribute to output error and selectively retains their FP16. Moreover, to ensure efficiency, we expose key intermediate variables from FlashAttention dao2023flashattention2 and integrate a lightweight GPU kernel for AnS computation during inference, enabling efficient runtime anchor token selection.
To validate AnTKV’s effectiveness, we conduct extensive experiments across various quantization bit-widths on a range of LLMs, including LLaMA-2/3 llama2 ; llama3 and Mistral-7B mistral family. As depicted in Figure 2, AnTKV consistently achieves better performance across a wide spectrum of bit-widths, from 4-bit down to 0.375-bit. On Wikitext2 with Mistral-7B, AnTKV achieves a perplexity of 6.32 at 1-bit, a reduction of 0.93 compared to CQ (7.25) zhang2024kv , and substantial 9.04 over KVQuant-1% (15.36) hooper2024kvquant . Even at an aggressive 0.375-bit, AnTKV further lowers perplexity to 8.87, surpassing CQ (13.20) by 4.33. Enabled by the efficient implementation of AnS, AnTKV scales to extremely long contexts. On LLaMA-3-8B, it supports up to 840K context length under 0.375-bit on a single A100-80GB GPU. During decoding, AnTKV increases the maximum batch size by 3.3x and improves throughput by 3.4x at 1K context length compared to FP16. More experiments that demonstrate the performance advantages of the proposed AnTKV are in Section 4.
To summarize, we make the following contributions in this work.
  • We conduct an error analysis of the self-attention, and introduce AnS to quantify the importance of tokens. With AnS, we can identify anchor tokens that significantly impact the accuracy of LLMs. Based on AnS, we present a novel token-aware VQ method, namely AnTKV, to compress the KV cache in LLM inference.
  • To ensure efficiency, we design a custom GPU kernel for computing AnS in the online stage that is compatible with FlashAttention. With the high-performance implementation of AnS, AnTKV can support up to 840K context length on a single GPU.
  • We deploy the proposed AnTKV in LLaMA-2/3 and Mistral-7B family, numerical results show that AnTKV has a significant performance improvement under various quantization bit-widths.
2 Background and Related Works
2.1 Transformer and Attention
The transformer block has become the fundamental architecture of LLMs; it consists of a fully connected feed-forward layer and a self-attention layer. The self-attention layer enables the model to capture connections among tokens within the context and serves as the critical component of the transformer block. Specifically, it maps a query ($\bm{Q}$) and a set of KV pairs ($\bm{K}$ and $\bm{V}$) to an output, i.e.,
$\texttt{Attn}(\bm{Q},\bm{K},\bm{V})=\texttt{Softmax}\left(\frac{\bm{Q}\bm{K}^{T}}{\sqrt{d}}\right)\bm{V}\quad\text{with}\quad\bm{Q,K,V}\in\mathbb{R}^{n\times d},$
where $\texttt{Softmax}(\cdot)$ is the softmax operator. Since position information is crucial in LLMs, Rotary Position Embedding (RoPE) su2023roformerenhancedtransformerrotary is a widely used technique for encoding it into query and key vectors su2024roformer , which is denoted as $\widetilde{\bm{Q}}$ and $\widetilde{\bm{K}}$ in this paper. With RoPE, the attention score $\bm{A}$ is rewritten as $\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}\right)$, and its $L_{p}$ “entry-wise” matrix norm is defined as $\left(\sum\limits_{i,j}|\bm{A}_{i,j}|^{p}\right)^{1/p}$. Other notations used in this work, such as the Kronecker product, the Hadamard product, and the row-by-row vectorization are denoted as $\otimes$, $\odot$, and $\texttt{Vec}(\cdot)$, respectively.
2.2 Memory Constraints in LLM Inference
During the process of LLM inference, the key and value generated at prefill and each decoding step are stored to avoid redundant computation. The KV cache is repeatedly accessed in subsequent steps to compute attention over the full context seen so far, significantly accelerating autoregressive decoding. In a multi-turn conversation, the KV cache from previous turns can also be reused during the prefill stage to further reduce latency. The state-of-the-art open-source and closed-source LLMs, like LLaMA-3.1-8B llama3 and Gemini 1.5 Pro team2023gemini , now handle much longer context lengths, up to 128K and 2 million tokens, respectively. This increase in context length liu2025comprehensive ; wang2025agenttaxo causes the GPU memory used for the KV cache in LLM inference to even exceed the models’ size. The rapid growth of KV cache greatly limits the deployment of LLMs with long context. Quantizing the KV cache directly optimizes this by maintaining key-value information and cutting memory costs substantially, thus enhancing the efficiency of LLM inference.
2.3 Related Works
KV cache quantization A variety of KV cache quantization methods have been developed to address the memory bottleneck in long-context LLMs liu2024kivi ; duanmu2024skvq ; hooper2024kvquant ; zhang2024kv ; zhu2024sampleattentionnearlosslessaccelerationlong . KIVI liu2024kivi reduces quantization error via per-channel key quantization combined with a sliding window strategy to emphasize locally relevant tokens. SKVQ duanmu2024skvq further explores this direction by introducing channel reordering and clipping to improve low-bit quantization. In contrast, KVQuant hooper2024kvquant adopts a different design by quantizing pre-RoPE keys using non-uniform quantization and selectively preserving a small fraction of full-precision KV entries to mitigate accuracy loss. CQ zhang2024kv adopts a VQ-based approach, aiming to exploit cross-channel correlations to further compress the KV cache.
KV cache compression Beyond quantization, the field of LLMs is actively exporing several advanced methods for KV cache compression. Sparse attention are designed to achieve compression by selectively computing attention over a reduced set of tokens xiaoefficient ; chen2024prefixquant ; zhu2025oraclekv ; liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; li2024snapkvllmknowslooking , However, a significant limitation lies in the inherent complexity of designing and optimizing their sparsity patterns. This often necessitates task-specific or data-driven heuristics, which can severely constrain their efficient generalization across diverse model architectures and datasets. Token merge, as a variant of sparse attention, merges the KV cache of similar tokens during inference to reduce memory usage zhang2024cam ; wang2024modeltellsmergeadaptive . Retrieval-based methods liu2024retrievalattention ; chenmagicpig ; zhang2024pqcache compress KV caches by retrieving relevant historical context for each query and are effective in tasks like RAG or document retrieval. However, they often lose substantial KV information, making them less suitable for general-purpose LLM inference.
Model Compression Many model compression techniques closely align with KV cache compression in both goals and underlying methodologies. GPTQ frantar2023gptqaccurateposttrainingquantization utilizes calibration set to reduce quantization induced degradation, while SmoothQuant xiao2024smoothquantaccurateefficientposttraining and AWQ lin2024awqactivationawareweightquantization minimize output error from the perspective of error propagation analysis. VQ-based methods such as VPTQ liu2024vptqextremelowbitvector and QUIP# tseng2024quipbetterllmquantization further enhance compression fidelity through residual VQ and Hadamard transform. Additionally, system-level frameworks like Atom zhao2024atomlowbitquantizationefficient and QServe lin2025qservew4a8kv4quantizationcodesign jointly optimize model and KV cache quantization, achieving end-to-end acceleration under low-bit quantization. Other works, such as Pruner-Zero dong2024prunerzeroevolvingsymbolicpruning and Parzc dong2024parzcparametriczerocostproxies , explore how to sparsify model weights while preserving model performance. Systems like FlashLLM xia2023flashllmenablingcosteffectivehighlyefficient and Spinfer fan2025spinfer leverage the model sparsity to accelerate inference.
3 Methodology
Previous works on quantifying the KV cache typically do not distinguish the importance of individual tokens, but assume that all tokens contribute equally to attention accuracy, or they adopt a simpler strategy of only preserving designated sink tokens. However, our 1-bit VQ experiment on Mistral-7B contradicts this, unveiling a previously unaddressed aspect in KV cache quantization: a small subset of tokens provokes disproportionately large perturbations in the self-attention outputs. This divergence reveals a new opportunity for KV cache quantization. Instead of uniform application, the precision of quantization should correspond to each token’s individual importance. Specifically, if we identify and preserve a subset of high-impact tokens in FP16, we can unlock a far better memory–accuracy trade-off.
Inspired by this, we present a novel token-aware VQ method for KV cache compression, namely, AnTKV. In AnTKV, we adopt a two-stage design: offline token-aware centroids learning (see Figure 3(a)) and online anchor token selection (see Figure 3(b)), the key of which lies in how to measure the importance of tokens. In the offline stage, we perform a weighted k-means clustering to produce centroids in VQ, and the gradient corresponding to each token is used as weights. In the online stage, to reduce the output error caused by quantization, we select tokens with both higher $\text{Err}_{\text{quant}}$ and AnS, i.e. anchor tokens, and keep them at FP16. Additionally, since applying RoPE disrupts the consistency of channel-wise magnitudes in $\bm{K}$ (see Appendix A for an illustrative example), which exhibits large inter-cluster distances and small intra-cluster variances, leading to large quantization error, the pre-RoPE strategy with the same as hooper2024kvquant is used in AnTKV.
Figure 3: The overview of AnTKV. (a) In the offline stage, we perform token-aware centroid learning to obtain the centroids. (b) In the online stage, we apply VQ using centroids and compute the AnS. Then identify the top-$k$ anchor tokens with the highest product of VQ error and AnS, and retain them in full-precision to preserve model performance.
3.1 Offline Token-Aware Centroids Learning
The offline stage aims to learn centroids by aggregating KV cache from general calibrations. Specifically, each row (per token) of $\bm{K}$ and $\bm{V}$ is divided into several sub-vectors, then using k-means to cluster them, and only the centroids are stored. To integrate the importance of tokens, we introduce the weighted k-means instead of classical k-means for the offline centroids learning.
From the perspective of accuracy, we hope that the errors caused by quantizing $\bm{K}$ and $\bm{V}$ will not affect the accuracy of LLMs as much as possible. Mathematically, it can be formulated as
$\begin{split}\min\limits_{\bm{C}^{k}\in\mathbb{R}^{c^{k}\times d}}|L(\bm{K},\bm{V})-L(\texttt{vq}(\bm{K}),\texttt{vq}(\bm{V}))|,\\ \end{split}$ (1)
where $c^{k}$ is the number of clusters, and $\bm{C}^{k}$ is the matrix consisting of centroids. By the first-order Taylor series expansion, we have
$\begin{split}L(\bm{K},\bm{V})-L(\texttt{vq}(\bm{K}),\texttt{vq}(\bm{V}))\approx\sum\limits_{j}\left(\langle\frac{\partial L}{\partial\bm{K}_{j,:}},\bm{K}_{j,:}-\texttt{vq}(\bm{K}_{j,:})\rangle+\langle\frac{\partial L}{\partial\bm{V}_{j,:}},\bm{V}_{j,:}-\texttt{vq}(\bm{V}_{j,:})\rangle\right),\end{split}$ (2)
where $\frac{\partial L}{\partial\bm{K}_{j,:}}$ and $\frac{\partial L}{\partial\bm{V}_{j,:}}$ are the gradient of the loss function $L(\cdot)$ with respect to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$, i.e., the $j$th row of $\bm{K}$ and $\bm{V}$ that correspond to the $j$th token. From (1) and (2), the weighted k-means clustering with gradients as weights is essentially to find a clustering strategy that minimizes the error caused by KV quantization. This also explains the effectiveness of the weighted k-means clustering used in this stage.
3.2 Online Anchor Token Selection
In the online stage, the gradient can no longer be used as AnS due to prohibitive memory and computational costs under real-time constraints. To this end, we here give the error propagation analysis of the self-attention (i.e., Attn). Specifically, the error propagation analysis is based on the perturbation bound of Attn that is related to each row of $\bm{K}$ and $\bm{V}$; see Theorem 3.1 (see Appendix B for proof).
Theorem 3.1.
Let $\delta\bm{K}$ and $\delta\bm{V}$ be the error perturbation terms corresponding to $\bm{K}$ and $\bm{V}$ respectively, and satisfy
$\|\delta\bm{K}\|_{L_{1}}\ll\|\bm{K}\|_{L_{1}}\quad\text{and}\quad\|\delta\bm{V}\|_{L_{1}}\ll\|\bm{V}\|_{L_{1}}.$
Then we have
$\|\texttt{Attn}(\bm{Q},\bm{K}+\delta\bm{K},\bm{V})-\texttt{Attn}(\bm{Q},\bm{K},\bm{V})\|_{L_{1}}$ (3)
$\lesssim$ $\sum_{j}\sum_{i}\bigl{\|}\bigl{(}\bm{V}^{\!\top}\mathrm{Diag}(\bm{A}_{i,:})(\bm{I}_{n}-\bm{e}\bm{A}_{i,:})\bigr{)}_{:,j}\bigr{\|}_{L_{1}}\|\bm{Q}_{i,:}\|_{L_{2}}\|\delta\bm{K}_{j,:}\|_{L_{1}}$
and
$\|\texttt{Attn}({\bm{Q}},{\bm{K}},\bm{V})-\texttt{Attn}({\bm{Q}},{\bm{K}},\bm{V}+\delta\bm{V})\|_{L_{1}}\leq\sum\limits_{j}\|\bm{A}_{:,j}\|_{L_{1}}\|\delta\bm{V}_{j,:}\|_{L_{1}},$ (4)
where $\bm{e}\in\mathbb{R}^{n}$ is a vector whose entries are all $1$.
We remark that the error propagation factors corresponding to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$ given in Theorem 3.1 can be regarded as the upper bound of the gradient of the self-attention operator related to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$.
The computation involving $\bm{K}$ in Theorem 3.1 entails substantial overhead, rendering it impractical for use during online inference. To address this, we propose a simplified variant that excludes the contribution of $\bm{V}$ to the quantization error of $\bm{K}$. By this simplification, we obtain the following reformulation. A detailed evaluation of the effectiveness of AnS is provided in Appendix C.
$\text{AnS}(\bm{K}_{j,:})=\sum_{i}\bm{A}_{i,j}(1-\bm{A}_{i,j})\cdot\left\|\bm{Q}_{i,:}\right\|_{2}$ (5)
During the prefill phase, AnS serves as an effective metric to identify anchor tokens that preserve critical contextual information by measuring their impact on attention outputs. In contrast, although AnS can still be computed during decoding, the anchor token it identifies may have already been quantized during prefill, preventing the preservation of its FP16 value for reducing accumulated errors. An important observation is that $\bm{A}$ and $\bm{A}\odot(\bm{E}-\bm{A})$ exhibit strong locality during decoding (see Appendix D), with anchor tokens primarily concentrated at the beginning and end of the sequence. In particular, the beginning tokens—referred to as sink tokens—are almost always identified as anchor tokens during prefill, indicating that our method effectively handle sink tokens. Motivated by this locality, we adopt a sliding window approximation of AnS during decoding.
3.3 Implementation
For the offline stage, the gradient of $\bm{K}$ and $\bm{V}$ is used as a weight to learn centroids. We implement it with a custom LinearWithAct to capture KV cache and corresponding gradients. Subsequently, we employ the weighted K-Means provided by cuML to perform efficient clustering.
For the online stage, AnS is derived from the error propagation factor given in Theorem 3.1 and (5). We design and implement a dedicated GPU kernel using Triton to efficiently compute it in conjunction with FlashAttention, enabling long-context inference. Since AnS requires reduction operations over the attention score matrix $\bm{A}$ and its transformed form $\bm{A}\odot(\bm{E}-\bm{A})$ along query (column-wise), directly fusing its computation into FlashAttention is impossible. To preserve the efficiency of FlashAttention, we decouple the AnS computation and execute it immediately afterward. To support this design, we extend FlashAttention to preserve three auxiliary tensors: the $L_{2}$ norm of each query vector, the key-wise (row-wise) sum, and the key-wise maximum of the matrix $\bm{Q}\bm{K}^{T}$. These additional outputs enable efficient reconstruction of the attention scores and allow AnS to be computed with minimal overhead. More details about the implementation are described in Algorithm 1 of Appendix E.
4 Experiments
In this section, we compare the performance of AnTKV with existing KV quantization methods through extensive experiments. The experimental setup, including models, benchmark datasets, evaluation metrics, and selected baselines, is as follows.
Models, Datasets, Metrics, and Parameter Settings. To validate the effectiveness of AnTKV in KV cache quantization and ensure the generality of results, we evaluate five representative models from the LLaMA and Mistral families: LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-8B-Instruct, Mistral-7B, and Mistral-7B-Instruct-v0.3. We use 128 samples of length 2048 from the WikiText2-v1 training set as the calibration. Perplexity is evaluated on WikiText-2 and C4, while zero-shot accuracy on MMLU hendrycks2021mmlu , ARC-C clark2018arc , MathQA amini2019mathqa , and PIQA bisk2020piqa is used to assess understanding and reasoning. Long-context performance is measured using LongBench bai2024longbench . For perplexity and zero-shot benchmarks, quantized KV cache is employed for attention. In LongBench, the FP16 KV cache is first utilized to compute attention outputs, before quantization is applied.
[…]
4.1 Perplexity Results
Perplexity is a standard metric that is widely used to evaluate the quality of the output of LLMs, with lower values indicating better performance. The perplexity results for different KV quantization approaches on WikiText-2 and C4 are presented in Table 1. The results in this table indicate that the proposed AnTKV consistently achieves competitive or superior perplexity across various bit-widths and model architectures. Under 4-bit and 2-bit quantization, AnTKV achieves competitive or superior performance compared to existing approaches. In the 1-bit and sub-bit regimes, AnTKV significantly outperforms all baselines. It is worth mentioning that on the C4 dataset under sub-bit quantization, the baseline methods exhibit extremely high perplexity due to anchor tokens not being well captured by fixed centroids. Thanks to effective AnS design and innovative Anchor Token selection, AnTKV substantially reduces perplexity, e.g., from 66.28 to 14.42 (LLaMA-3-8B, 0.75-bit).
Table 1: The evaluations are conducted using each model’s maximum context length, specifically 4096 for LLaMA-2-7B and 8192 for LLaMA-3-8B and Mistral-7B. "Ours" refers to our method (AnTKV) without using any anchor tokens, while Ours-1% denotes AnTKV where 1% of the tokens are treated as anchor tokens and retained in FP16. For simplicity, the reported bit-widths do not consider the effect of centroids.
Bit LLaMA-2-7B LLaMA-3-8B Mistral-7B
Dataset WikiText2 C4 WikiText2 C4 WikiText2 C4
Baseline 16 5.12 6.63 5.54 7.10 4.73 5.66
RTN 4 5.66 7.31 7.89 8.79 7.34 5.91
SKVQ 5.16 6.67 5.64 7.19 4.97 5.68
KVQuant-1% 5.13 6.65 5.56 7.12 4.78 5.72
CQ 5.14 6.67 5.58 7.84 4.79 5.74
Ours 5.18 6.76 5.61 7.69 4.76 5.69
Ours-1% 5.15 6.68 5.59 7.16 4.74 5.67
RTN 2 4708 4708 2841 2113 573 477
SKVQ 5.54 7.21 6.73 8.31 5.21 6.14
KVQuant-1% 5.49 7.02 6.11 7.65 5.19 6.10
CQ 5.42 7.23 6.09 18.71 5.11 6.17
Ours 5.51 7.45 6.10 16.96 5.08 6.18
Ours-1% 5.34 7.02 5.97 7.68 4.95 5.97
SKVQ 1 12643 12819 108879 86426 3524 2741
KVQuant-1% 21.55 51.84 14.80 13.95 15.36 14.24
CQ 7.75 12.49 9.56 81.74 7.25 9.89
Ours 7.92 13.01 9.62 74.47 7.32 10.51
Ours-1% 6.50 9.40 8.51 12.51 6.32 8.44
CQ 0.75 8.39 14.32 11.18 72.05 7.64 11.72
Ours 8.21 14.27 10.41 66.28 7.41 11.72
Ours-1% 6.55 9.75 8.97 14.42 6.43 9.08
CQ 0.375 14.82 33.59 22.80 103.5 13.20 26.34
Ours 13.37 30.51 17.70 103.5 11.65 23.98
Ours-1% 8.75 15.86 13.41 34.08 8.87 14.87
4.2 Understanding and Reasoning Benchmark
To ensure that AnTKV exhibits broad understanding and reasoning capabilities, we evaluate its performance on four representative benchmarks: MMLU, ARC-Challenge, PIQA, and MathQA, using both LLaMA-8B-Instruct and Mistral-7B-Instruct. These benchmarks respectively assess multi-domain knowledge reasoning (MMLU), complex question answering (arc_challenge), commonsense reasoning (PIQA), and mathematical problem-solving (MathQA). Figure 4 illustrates the comparison results of AnTKV with SKVQ and CQ. Similarly, it can be seen that AnTKV consistently outperforms SKVQ and CQ across all four benchmarks under low-bit quantization, especially at 1-bit and sub-bit levels.
Figure 4: Understanding and Reasoning ability evaluation accuracy results on MMLU, ARC-C, PIQA, and MathQA under different quantization bit-widths.
4.3 Long-Context Benchmark
To validate the effectiveness of AnTKV in handling long-context, We conduct several experiments on the LLaMA-8B-Instruct model using the LongBench benchmark, a diverse collection of tasks such as question answering, retrieval, and summarization, designed to systematically evaluate long-context understanding in language models. We report results on eleven representative sub-tasks from LongBench, along with averaged performance. Due to alignment issues of KIVI and SKVQ, we exclude the triviaqa and gov_report sub-tasks from the comparison. As shown in Figure 5, AnTKV preserves nearly FP16 at 4- and 2-bit quantization across almost all tasks. At the 1-bit quantization, the performance of KIVI and SKVQ has a significant drop. In contrast, AnTKV and CQ still maintain a relatively high accuracy. To further investigate the robustness under aggressive compression, we compare AnTKV and CQ in both sub-bit levels. Figure 5 shows that AnTKV consistently outperforms CQ. Notably, when quantizing down to 0.375-bit, AnTKV maintains a tolerable performance degradation, with the average score dropping from 46.2 to 34.5.
Figure 5: The evaluation accuracy results on LongBench under different KV cache quantization bit-widths.
4.4 Efficiency
In this experiment, we evaluate the efficiency of our AnTKV implementation against the FlashAttention-accelerated FP16 baseline on LLaMA-3-8B using a single A100-80GB GPU. As shown in Figure 6, AnTKV substantially extends the maximum context length from 128K to 384K. In long-context scenarios, we observe that intermediate activations consume a significant portion of memory. By applying a series of in-place operations, AnTKV ultimately supports up to 810K tokens under 1-bit quantization and 840K under 0.375-bit quantization. More importantly, AnTKV sustains low memory consumption while maintaining acceptable latency. Under 1-bit quantization, a noticeable increase in prefill latency is observed only at the 128K context length, which represents the upper limit supported by most current models. It is observed that 0.375-bit quantization introduces higher latency than its 1-bit. This increase in latency arises from the use of a larger number of centroids, which are employed to enhance model accuracy.
To evaluate decoding efficiency, we evaluate the throughput of AnTKV under a fixed context length of 1K tokens. As shown in Figure 7, AnTKV supports significantly larger batch sizes compared to the baseline method. By substantially reducing KV cache access, it improves throughput across all bit-widths. Particularly, the maximum throughput reaches 3.5x that of the baseline method under 1-bit quantization.
Figure 6: Memory and latency comparison. Bars show memory usage (red: 1-bit, green: 0.375-bit); lines indicate prefill latency across context lengths.
Figure 7: Decoding throughput comparison. Our method supports larger batch sizes and achieves higher throughput.
4.5 Ablation Study
We conduct a series of experiments to answer the following questions.
Q1: How does the model performance change as the number of anchor tokens increases?
As the number of anchor tokens increases, the performance loss decreases rapidly at first, as shown in Table 1. However, the marginal benefit diminishes with more anchor tokens. Detailed results are provided in Appendix F.
Q2: Does the calibration set affect the performance?
We find that for VQ-based methods, the calibration set does have some impact on performance under low-bit settings. However, as shown in Table 1 (LLaMA-3-8B, 1-bit, C4), retaining anchor tokens effectively mitigates the performance drop caused by calibration set variation. More detailed results can be found in Appendix G.
5 Limitation & Conclusion
Although AnTKV demonstrates its advantages in experiments, it also has several limitations. First, more accurate AnS for tokens and higher performance implementations for its computation may be possible. AnTKV demonstrates strong potential in LLM serving by substantially reducing the size of the KV cache, which in turn alleviates I/O and memory constraints to a significant extent. Nevertheless, further empirical validation is required.
We observed from several experiments that different tokens exhibit varying sensitivity to KV quantization, which opens a new opportunity for a better memory-accuracy trade-off. To this end, we present AnTKV, a two-stage token-aware VQ framework that involves offline gradient weighted centroid learning and online anchor token selection guided by AnS. Moreover, to ensure the efficiency of AnTKV in LLM inference, we implement a lightweight GPU kernel compatible with FlashAttention to compute AnS with minimal overhead. Extensive experiments show that AnTKV achieves superior performance under low and ultra-low-bit quantization, significantly extending the maximum context length (up to 840K tokens) on a single A100 GPU while maintaining low perplexity and high throughput. Our results demonstrate the potential of token-aware quantization in scaling LLM inference to longer contexts without sacrificing accuracy.
Appendix A Distribution of pre- and post- RoPE Key
To identify a quantization strategy better suited for $\bm{K}$ vertor quantization, we compare the distribution of $\bm{K}$ before and after applying RoPE. Figure 8 presents a visualization of the pre- and post- RoPE $\bm{K}$. We observe that, compared to the post-RoPE $\bm{K}$, the pre-RoPE $\bm{K}$ exhibits smaller inter-cluster distances and lower intra-cluster variance, which contributes to reduced quantization error.
Figure 8: Distribution of pre- and post- RoPE Key. We sampled a 2048-length sentence from WikiText2 and generated pre- and post- RoPE Key on the LLaMA-3-8B model.
Appendix B Proof of Theorem 3.1
For $\bm{K}$, we have
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K}+\delta\bm{K},\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}=\|\left(\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}\widetilde{\delta\bm{K}}^{T}}{\sqrt{d}}\right)-\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}^{T}\widetilde{\bm{K}}}{\sqrt{d}}\right)\right)\bm{V}\|_{L_{1}}.\end{split}$ (6)
The key to estimating the bound of (6) lies in the analysis of
$\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}\widetilde{\delta\bm{K}}^{T}}{\sqrt{d}}\right)-\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}\right),$ (7)
whose $(i,j)$th entry is represented as
$\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}-\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}.$ (8)
Since $\|\delta\bm{K}\|_{L_{1}}\ll\|\bm{K}\|_{L_{1}}$, and by the first-order approximation $\exp{(x+\delta x)}\approx\exp{(x)}(1+\delta x)$, (8) can be approximated as
$\begin{split}\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\right)-\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)\right)}{\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\right)\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)\right)}\\ =\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}\cdot\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\\ \approx\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}\cdot\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}}\\ =\bm{A}_{i,j}\cdot\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}}.\end{split}$
For convenience, we denote $\frac{\widetilde{\bm{Q}}\widetilde{\delta\bm{K}}^{T}}{\sqrt{d}}$ and $\left[\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}}\right]_{n\times n}$ as $\bm{X}$ and $\bm{Y}$ respectively. Then (7) can be approximated as $\left(\bm{A}\odot\bm{Y}\right)\bm{V}$, and by the property of Kronecker product, we have
$\begin{split}\texttt{Vec}\left((\bm{A}\odot\bm{Y})\bm{V}\right)=\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Vec}\left(\bm{A}\odot\bm{Y}\right)=\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Diag}\left(\texttt{Vec}(\bm{A})\right)\texttt{Vec}(\bm{Y}).\end{split}$
Further, we can obtain
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K}+\delta\bm{K},\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}&\approx\|\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Diag}\left(\texttt{Vec}(\bm{A})\right)\texttt{Vec}(\bm{Y})\|_{L_{1}}\\ &=\sum\limits_{i}\|\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\bm{X}_{i,:}^{T}\|_{L_{1}}\\ &\leq\sum\limits_{i}\sum\limits_{j}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}|\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}|\\ &\leq\sum\limits_{j}\sum\limits_{i}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}\|{\bm{Q}}_{i,:}\|_{2}\|{\delta\bm{K}}_{j,:}\|_{L_{1}}.\\ \end{split}$ (9)
For $\bm{V}$, we have
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}+\delta\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}&=\|\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}\right)\delta\bm{V}\|_{L_{1}}\\ &=\|\bm{A}\delta\bm{V}\|_{L_{1}}=\sum\limits_{i,k}|\sum\limits_{j}\bm{A}_{i,j}\delta\bm{V}_{j,k}|\\ &\leq\sum\limits_{i,k}\sum\limits_{j}\bm{A}_{i,j}|\delta\bm{V}_{j,k}|\\ &=\sum\limits_{j}\left(\sum\limits_{i}\bm{A}_{i,j}\right)\left(\sum\limits_{k}|\delta\bm{V}_{j,k}|\right)\\ &=\sum\limits_{j}\|\bm{A}_{:,j}\|_{L_{1}}\|\delta\bm{V}_{j,:}\|_{L_{1}}.\end{split}$ (10)
Appendix C The Effectiveness of AnS
We quantized each token across layers and heads, and separately recorded the errors in the attention outputs. As shown in Figure 9 and (5), it can be observed that the relative values derived from our AnS, as defined in Theorem 3.1, closely align with the actual outputs of the attention. For a fair comparison, the output errors for $\bm{K}$ also exclude the contribution of $\bm{V}$.
Figure 9: The effectiveness of AnS.
Appendix D AnS Distribution During Decoding
To illustrate the distribution of AnS during decoding, we sampled prompts from Qasper within LongBench for visualization. As shown in Figure 10, we present the distribution of AnS($\bm{K}$) and AnS($\bm{V}$) across different layers and heads, specifically when decoding the first token. Our observations reveal that high AnS values during decoding are predominantly concentrated on adjacent tokens and at the attention sink tokens. Since sink tokens often lead to significant error propagation and can be dynamically identified by AnS during prefill, we simplify the design of AnS during decoding by employing a sliding window to ensure model performance.
Figure 10: AnS Distribution on Sampled Prompts from Qasper Using LLaMA-3-8B-Instruct During First-Token Decoding.
Appendix E Computational Procedure of AnS in the Online Stage
Algorithm 1 The computation of AnS in the online stage.
1:  Input: Query ($\bm{Q}$), key ($\bm{K}$), value ($\bm{V}$)
2:  Output: AnS of KV, i.e., $\text{AnS}({\bm{K}})$ and $\text{AnS}({\bm{V}})$
3:  $(\bm{O},\bm{L},\bm{M},\left\|\bm{Q}_{i,:}\right\|_{L_{2}})\leftarrow\texttt{FlashAttention}(\bm{Q},\bm{K},\bm{V})$
4:  for each block key index $j$ in parallel (assigned to GPU block) do
5:     for each block query index $i$ do
6:        $\bm{S}_{i,j}\leftarrow\langle\bm{Q}_{h,i},\bm{K}_{h,j}\rangle$
7:        $\bm{A}_{i,j}\leftarrow\exp(\bm{S}_{i,j}-\bm{M}_{h,i})/\bm{L}_{h,i}$
[…]
10:     end for
11:  end for
12:  return $\text{AnS}({\bm{K}})$, $\text{AnS}({\bm{V}})$
Appendix F Calibration Set Impact
As shown in the Table 2, we observe that for VQ-based quantization in the ultra-low-bit regime, the calibration set significantly impacts the perplexity results. However, AnTKV with $1\%$ anchor tokens not only substantially reduces the PPL but also greatly mitigates the effect of different calibration sets.
Table 2: Perplexity experiment results on Mistral-7B, using the W2 and C4 training sets respectively as Calibration Sets. The "Vset" is the validation set related to W2 and C4. "Calib Set" represents "Calibration Set".
Bits Vset 4 2 1 0.75 0.375
Calib Set W2 C4 W2 C4 W2 C4 W2 C4 W2 C4
Ours W2 4.76 5.69 5.08 6.18 7.32 10.51 7.32 10.51 11.65 23.98
C4 4.79 5.69 5.32 6.15 10.14 10.09 10.90 10.80 24.16 19.95
Ours-1% W2 4.74 5.67 4.95 5.97 6.32 8.44 6.32 8.44 8.87 14.87
C4 4.75 5.66 5.02 5.94 7.13 8.13 7.79 8.56 12.87 13.07
To further investigate the impact of the calibration set on model performance, we used C4 as a calibration set to evaluate several subtasks within LongBench (qasper, trec, samsum, lcc, ropebench-p). As shown in the Table 3, we observed that there were some differences in the results of Trec and Repobench-p when using Wikitext-2 and C4, while the differences were not significant for the other tasks.
Table 3: Performance on LongBench Subtasks with Wikitext-2 (W2) and C4 Calibration Sets at Different Bits using LLaMA-3-8B-Instruct. "Calib Set" represents "Calibration Set", and "repobc-p" represents "repobench-p".
Bits 4 2 1 0.75 0.375
Calib Set W2 C4 W2 C4 W2 C4 W2 C4 W2 C4
qasper 40.46 39.98 39.04 38.2 25.95 26.51 25.48 25.49 22.41 23.27
trec 69.33 69.33 67 64.67 38.67 42.33 39.67 41.33 38 38
samsum 40.2 40.27 38.61 38.22 30.0 30.3 29.57 29.29 25.5 24.82
lcc 59.84 59.07 60.94 59.15 53.97 53,93 52.97 52.01 49.61 49.79
repobc-p 44.24 41.3 45.29 42.68 38.53 37.94 37.87 38.02 34.54 34.71
Appendix G Anchor Tokens Number Impact
To investigate the impact of the number of anchor tokens on model performance, we conducted Perplexity evaluations on Mistral-7B and LLaMA-3-8B, both with a context length of 8192. We performed evaluations using no anchor tokens and with anchor token percentages of 1% (82), 2% (164), 5% (410), 10% (820), 15% (1230), and 20% (1640). The corresponding results are presented in Figures 11 and 12. For the 2-bit and 4-bit results, using $1\%$ of anchor tokens kept the error within 0.6 compared to FP16. However, for the 1-bit and sub-bit results, we needed to increase the number of anchor tokens to control the error within an acceptable range. Nevertheless, AnTKV provides a feasible technical pathway for ultra-low-bit quantization of the KV cache.
Figure 11: Perplexity results on LLaMA-3-8B with varying anchor token numbers.
Figure 12: Perplexity results on Mistral-7B with varying anchor token numbers.
To further investigate the impact of anchor token numbers on downstream tasks, we evaluated different anchor token numbers on the Trec and Qasper subtasks of LongBench under ultra-low-bit quantization settings. For convenience, we approximated $1\%$ of the anchor token number as 64. The results are shown in Figure 13. The results indicate that Trec’s performance increases linearly with the number of anchor tokens, eventually reaching a result within 10 points of the baseline when using $10\%$ anchor tokens. In contrast to Trec, Qasper’s accuracy remains stable with a small number of anchor tokens but shows a significant performance improvement when increased to $10\%$. We believe this difference stems from the inherent properties of each task.
Figure 13: Trec and Qasper results on LLaMA-3-8B-Instruct with varying anchor token numbers.
Appendix H Models, Datasets and Code
H.1 Models Checkpoints
H.2 Datasets
H.3 Code
Our code and the trained centroids will be publicly released upon the acceptance of this paper.
Appendix I Broader Impact
In this work, we observe that tokens in the KV cache exhibit heterogeneous sensitivity to quantization, an insight that challenges the prevailing assumption of uniform importance across tokens. This observation paves the way for a new paradigm of importance-aware compression in LLMs. Our proposed method, AnTKV, embodies this paradigm through a two-stage, token-aware vector quantization framework. By selectively preserving high-impact tokens, AnTKV achieves strong performance even under ultra-low-bit quantization. Given its effectiveness and compatibility with high-throughput inference, AnTKV holds the potential to be widely adopted in future LLM compression and deployment practices.

### v2

@@
 Abstract
-Quantization has emerged as an effective and lightweight solution to reduce the memory footprint of the KV cache in Large Language Models (LLMs).
-Nevertheless, minimizing the performance degradation caused by ultra-low-bit KV cache quantization remains a significant challenge.
-We observe that quantizing the KV cache of different tokens has varying impacts on the quality of attention outputs.
-To systematically investigate this phenomenon, we perform forward error propagation analysis on attention and propose the Anchor Score (AnS) that quantifies the sensitivity of each token’s KV cache to quantization-induced error.
-Our analysis reveals significant disparities in AnS across tokens, suggesting that preserving a small subset with full precision (FP16) of high-AnS tokens can greatly mitigate accuracy loss in aggressive quantization scenarios.
-Based on this insight, we introduce AnTKV, a novel framework that leverages Anchor Token-aware Vector Quantization to compress the KV cache.
-Furthermore, to support efficient deployment, we design and develop a triton kernel that is fully compatible with FlashAttention, enabling fast online Anchor Token selection.
-AnTKV enables LLaMA-3-8B to handle context lengths up to 840K tokens on a single 80GB A100 GPU, while achieving up to 3.5x higher decoding throughput compared to the FP16 baseline.
-Our experiment results demonstrate that AnTKV matches or outperforms prior works such as KIVI, SKVQ, KVQuant, and CQ under 4-bit settings.
-More importantly, AnTKV achieves significantly lower perplexity under ultra-low-bit quantization on Mistral-7B, with only 6.32 at 1-bit and 8.87 at 0.375-bit, compared to the FP16 baseline of 4.73.
+Quantization has emerged as an effective and lightweight solution to reduce the memory footprint of the KV cache in Large Language Models.
+Nevertheless, minimizing the accuracy degradation caused by ultra-low-bit KV cache quantization remains a significant challenge.
+While scalar quantization is constrained by 1-bit bound, vector quantization exploits intra-vector correlations and enables sub-bit regimes, making it more suitable for ultra-low-bit quantization.
+To further mitigate quantization-induced degradation, we reveal that the degradation is highly uneven across tokens in attention quality.
+To investigate this unevenness, we introduce anchor score to measure each token’s sensitivity to quantization.
+Our analysis and experiments show that preserving a small subset (1%) of tokens with the highest Anchor Score significantly mitigates accuracy loss under aggressive quantization.
+We propose AnTKV, a dual-stage framework that leverages anchor token-aware vector quantization to compress the KV cache.
+It combines offline token-aware centroids learning and online anchor token selection to balance compression and accuracy.
+To enable efficient deployment, we design an online anchor token selection kernel compatible with FlashAttention.
+It allows LLaMA3-8B to scale to 840K tokens on a single 80GB A100, while delivering up to $3.5\times$ higher decoding throughput over the FP16 baseline.
+Experiments demonstrate that AnTKVmatches or surpasses prior methods at 4-bit, and significantly reduce perplexity under ultra-low-bit quantization, achieving 6.32 at 1-bit on Mistral-7B, compared to 7.25 for CQ and 15.36 for KVQuant.
 […]
 1 Introduction
-LLMs have attracted much attention due to their strong capability in natural language processing and other artificial intelligence tasks openai2023gpt4 ; guo2025deepseek ; team2023gemini ; tang2025the ; dong2025can .
-With the rapid development in recent years, LLM can manage increasingly larger context lengths ranging from hundreds of thousands to millions of tokens, and its broadening scope allows LLM to tackle more complex tasks zhao2023survey ; wang2025all ; das2025security ; wang2025agenttaxo .
-However, these models, which are typically built on decoder-based transformer architectures, generate tokens recursively during inference, thus leading to a large accumulation of the KV cache as the context length zhu2025oraclekv ; bai2024longbench ; liu2025comprehensive and batch size increases pope2022efficientlyscalingtransformerinference .
-The large memory footprint of the KV cache presents a significant challenge.
-For long-context inference, such as LLaMA-3’s 128K context length, the KV cache can consume several gigabytes of memory, nearly matching the model’s weights.
-This challenge is even greater for models like Gemini, which handle contexts up to a million tokens.
-Moreover, LLM-RAG systems yao2025cacheblend pre-generate and store KV caches.
-However, the excessive size of these caches presents a significant challenge for storage.
-Figure 1: Self-attention output error when quantizing the KV cache of Mistral-7B to 1-bit.
-Figure 2: The perplexity of Mistral-7B on the Wikitext2 under different bit-widths.
-To address the growing KV cache, various optimization strategies have been proposed liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; liu2024retrievalattention ; zhang2024cam ; liu2024kivi ; dong2025can , with quantization emerging as a prominent solution.
-Quantization approaches fall into two main categories: scalar quantization (SQ) and vector quantization (VQ).
-SQ compresses the KV cache by mapping floating-point values to fixed low-bit representations liu2024kivi ; duanmu2024skvq ; hooper2024kvquant , but the minimum 1-bit requirement limits its compression to 1/16 on FP16.
-In contrast, VQ maps high-dimensional vectors to a finite codebook lingle2024transformervq ; zhang2024kv , capturing intra-vector correlations.
-This enables higher compression rates and better preserves representational fidelity, especially under ultra-low-bit regimes.
-During the LLM inference, different tokens have different effects on the accuracy of LLMs, so it is unreasonable to treat all tokens equally in VQ methods.
-To illustrate this, we quantize the KV cache of Mistral-7B mistral to 1-bit using VQ, and Figure 1 shows the impact of KV quantization on the self-attention output error within layer 31.
-It reveals that even in the absence of tokens with exceptionally large quantization errors, a few tokens (Anchor Tokens) can still introduce substantial deviations in the output of the self-attention operator.
-Therefore, measuring the importance of tokens during inference can further explore the performance potential of VQ methods for compressing the KV cache.
-To this end, we introduce Anchor Score (AnS), a metric derived from forward error analysis, to quantify each token’s impact on the output.
-Leveraging AnS, we propose AnTKV, a novel framework that employs anchor token-aware VQ for KV cache compression.
-AnTKV mitigates quantization-induced degradation through a two-stage design: offline token-aware centroids learning and online anchor token selection.
-In the offline stage, the framework performs the token-aware weighted k-means clustering to generate centroids.
-During the online stage, it computes the AnS for each request to identify anchor tokens that significantly contribute to output error and selectively retains their FP16.
-Moreover, to ensure efficiency, we expose key intermediate variables from FlashAttention dao2023flashattention2 and integrate a lightweight GPU kernel for AnS computation during inference, enabling efficient runtime anchor token selection.
-To validate AnTKV’s effectiveness, we conduct extensive experiments across various quantization bit-widths on a range of LLMs, including LLaMA-2/3 llama2 ; llama3 and Mistral-7B mistral family.
-As depicted in Figure 2, AnTKV consistently achieves better performance across a wide spectrum of bit-widths, from 4-bit down to 0.375-bit.
-On Wikitext2 with Mistral-7B, AnTKV achieves a perplexity of 6.32 at 1-bit, a reduction of 0.93 compared to CQ (7.25) zhang2024kv , and substantial 9.04 over KVQuant-1% (15.36) hooper2024kvquant .
-Even at an aggressive 0.375-bit, AnTKV further lowers perplexity to 8.87, surpassing CQ (13.20) by 4.33.
-Enabled by the efficient implementation of AnS, AnTKV scales to extremely long contexts.
-On LLaMA-3-8B, it supports up to 840K context length under 0.375-bit on a single A100-80GB GPU.
-During decoding, AnTKV increases the maximum batch size by 3.3x and improves throughput by 3.4x at 1K context length compared to FP16.
-More experiments that demonstrate the performance advantages of the proposed AnTKV are in Section 4.
+Large Language Models (LLMs) have gained wide attention owing to their remarkable capabilities in diverse applications openai2023gpt4 ; guo2025deepseek ; team2023gemini ; tang2025the ; dong2025can .
+With rapid recent advances, LLMs currently handle context lengths from hundreds of thousands to millions of tokens, enabling them to tackle increasingly complex tasks zhao2023survey ; wang2025all ; das2025security ; wang2025agenttaxo .
+Most LLMs adopt decoder-based transformer architectures, where tokens are generated autoregressively and the KV cache grows rapidly with context length zhu2025oraclekv ; bai2024longbench ; liu2025comprehensive and batch size increases pope2022efficientlyscalingtransformerinference .
+The large amount of memory footprint of the KV cache poses a significant challenge.
+For example, with LLaMA-3 at 128K tokens it already approaches the model size, and for million-token models like Gemini it becomes prohibitive.
+Beyond inference, LLM-RAG systems yao2025cacheblend pre-generate and store massive KV caches, creating additional storage challenges.
+Figure 1: The $L_{1}$ norm error of attention output when quantizing the $i$th token’s KV cache in Mistral-7B to 1-bit.
+Figure 2: The Perplexity of Mistral-7B on the WikiText-2 across different quantization bit-widths.
+To address the substantial size of KV caches, numerous techniques have been explored liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; liu2024retrievalattention ; zhang2024cam ; liu2024kivi ; dong2025can , with quantization proving to be especially effective.
+Quantization methods are generally classified into scalar quantization (SQ) and vector quantization (VQ).
+SQ compresses the KV cache by mapping floating-point values to fixed low-bit representations liu2024kivi ; duanmu2024skvq ; hooper2024kvquant , but the 1-bit lower bound constrains its maximum compression ratio to 1/16 of FP16.
+In contrast, VQ compresses high-dimensional vectors by mapping them to a finite codebook lingle2024transformervq ; zhang2024kv , which captures intra-vector correlations, achieves superior compression ratios and fidelity under ultra-low-bit quantization.
+We observe that tokens contribute unequally to model accuracy during inference.
+To illustrate this, we quantize the KV cache of Mistral-7B mistral to 1-bit using VQ.
+Figure 1 depicts the attention output $L_{1}$ norm error in the 31st layer when quantizing the KV cache across different token indices, and similar error distributions are observed in other layers.
+As evident from the figure, although per-token quantization errors are similar, the resulting error in the attention output vary widely across tokens and a small subset exhibits error that are tens of times larger when quantized (anchor tokens).
+Quantifying token importance is essential for unlocking the full potential of ultra-low-bit KV cache quantization.
+To address this, we conduct a forward error analysis forwarderroranalysis on attention-based models with respect to KV, and then propose AnTKV, a dual-stage framework for KV cache quantization.
+In the offline stage, AnTKV performs token-aware weighted k-means clustering to generate centroids, where the weights are derived from error propagation factors obtained through forward error analysis.
+Tokens that cause larger increases in output error are assigned larger weights during clustering.
+However, the propagation factors entail substantial gradient cost.
+To overcome this limitation, we propose Anchor Score (AnS) derived from the forward error analysis of the attention, which quantifies the output sensitivity to quantizing each token’s KV cache.
+At inference time, AnS is computed for each prompt to identify this small token subset.
+AnTKV handles this subset of tokens in a simple yet effective manner by preserving them in full precision, thereby reducing accuracy loss.
+Moreover, we design and implement a lightweight GPU kernel for AnS computation.
+More specifically, we extend FlashAttention dao2023flashattention2 to store low memory overhead softmax intermediate results, thereby enabling efficient online anchor token selection.
+We conduct extensive experiments to evaluate the effectiveness of AnTKV across various quantization bit-widths on a range of LLMs, including the LLaMA-2/3 llama2 ; llama3 and Mistral-7B mistral families.
+As shown in Figure 2AnTKV consistently outperforms existing approaches across bit-widths from 4-bit down to 0.375-bit.
+On WikiText-2 wikitext2 AnTKV achieves a perplexity of 6.32 at 1-bit, reducing error by 0.93 compared to CQ (7.25) zhang2024kv , and by a substantial 9.04 compared to KVQuant-1% (15.36) hooper2024kvquant .
+Even under the aggressive 0.375-bit setting, AnTKV attains a perplexity of 8.87, surpassing CQ (13.20) by 4.33.
+Thanks to the efficient GPU implementation of AnS, AnTKV also scales to extremely long contexts: on LLaMA3-8B, it supports up to 840K tokens under 0.375-bit quantization on a single A100-80GB GPU.
+Moreover, during decoding, AnTKV increases the maximum batch size by $3.3\times$ and improves throughput by $3.4\times$ at a 1K context length compared to full precision.
 To summarize, we make the following contributions in this work.
-  • We conduct an error analysis of the self-attention, and introduce AnS to quantify the importance of tokens.
-With AnS, we can identify anchor tokens that significantly impact the accuracy of LLMs.
-Based on AnS, we present a novel token-aware VQ method, namely AnTKV, to compress the KV cache in LLM inference.
-  • To ensure efficiency, we design a custom GPU kernel for computing AnS in the online stage that is compatible with FlashAttention.
-With the high-performance implementation of AnS, AnTKV can support up to 840K context length on a single GPU.
-  • We deploy the proposed AnTKV in LLaMA-2/3 and Mistral-7B family, numerical results show that AnTKV has a significant performance improvement under various quantization bit-widths.
+  • To the best of our knowledge, this work is the first to investigate the feasibility of quantization the KV cache to sub-bit while preserving model accuracy.
+  • We identify that different tokens contribute unequally to model accuracy under quantization and highlight the existence of anchor tokens that dominate output error.
+  • We propose AnTKV, which performs token-aware weighted clustering offline and leverages AnS online to efficiently identify anchor tokens, preserving them in full precision to mitigate accuracy loss, and we evaluate it on the LLaMA2/3 and Mistral families, where it consistently improves performance across a wide range of quantization bit-widths.
+  • We implement custom GPU kernels for the online stage, enabling AnTKV to scale to 840K tokens on a single GPU and deliver significant throughput gains during decoding.
 2 Background and Related Works
 2.1 Transformer and Attention
-The transformer block has become the fundamental architecture of LLMs; it consists of a fully connected feed-forward layer and a self-attention layer.
-The self-attention layer enables the model to capture connections among tokens within the context and serves as the critical component of the transformer block.
+The transformer block has become the fundamental architecture of LLMs; it consists of a fully connected feed-forward network and an attention.
+The attention enables the model to capture connections among tokens within the context.
 Specifically, it maps a query ($\bm{Q}$) and a set of KV pairs ($\bm{K}$ and $\bm{V}$) to an output, i.e.,
@@
 In a multi-turn conversation, the KV cache from previous turns can also be reused during the prefill stage to further reduce latency.
-The state-of-the-art open-source and closed-source LLMs, like LLaMA-3.1-8B llama3 and Gemini 1.5 Pro team2023gemini , now handle much longer context lengths, up to 128K and 2 million tokens, respectively.
-This increase in context length liu2025comprehensive ; wang2025agenttaxo causes the GPU memory used for the KV cache in LLM inference to even exceed the models’ size.
+Nowadays LLMs, like LLaMA3.1-8B llama3 and Gemini 1.5 Pro team2023gemini , now handle much longer context lengths, up to 128K and 2 million tokens, respectively.
+This increase in context length liu2025comprehensive ; wang2025agenttaxo causes the GPU memory consumed by KV cache to even exceed the model.
 The rapid growth of KV cache greatly limits the deployment of LLMs with long context.
-Quantizing the KV cache directly optimizes this by maintaining key-value information and cutting memory costs substantially, thus enhancing the efficiency of LLM inference.
+Quantization is an effective approach to compressing the KV cache, substantially reducing memory footprint while preserving essential information.
 2.3 Related Works
-KV cache quantization A variety of KV cache quantization methods have been developed to address the memory bottleneck in long-context LLMs liu2024kivi ; duanmu2024skvq ; hooper2024kvquant ; zhang2024kv ; zhu2024sampleattentionnearlosslessaccelerationlong .
-KIVI liu2024kivi reduces quantization error via per-channel key quantization combined with a sliding window strategy to emphasize locally relevant tokens.
-SKVQ duanmu2024skvq further explores this direction by introducing channel reordering and clipping to improve low-bit quantization.
-In contrast, KVQuant hooper2024kvquant adopts a different design by quantizing pre-RoPE keys using non-uniform quantization and selectively preserving a small fraction of full-precision KV entries to mitigate accuracy loss.
+KV cache quantization A variety of KV cache quantization methods have been proposed to address the memory bottleneck in long-context LLMs liu2024kivi ; duanmu2024skvq ; hooper2024kvquant ; zhang2024kv ; zhu2024sampleattentionnearlosslessaccelerationlong .
+KIVI liu2024kivi mitigates quantization error by applying per-channel key quantization and employing a sliding window to emphasize locally relevant tokens.
+SKVQ duanmu2024skvq further explores this direction by introducing channel reordering and clipping.
+To further reduce accuracy loss, KVQuant hooper2024kvquant introduces pre-RoPE key quantization, non-uniform format and element-wise outlier.
 CQ zhang2024kv adopts a VQ-based approach, aiming to exploit cross-channel correlations to further compress the KV cache.
-KV cache compression Beyond quantization, the field of LLMs is actively exporing several advanced methods for KV cache compression.
-Sparse attention are designed to achieve compression by selectively computing attention over a reduced set of tokens xiaoefficient ; chen2024prefixquant ; zhu2025oraclekv ; liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; li2024snapkvllmknowslooking , However, a significant limitation lies in the inherent complexity of designing and optimizing their sparsity patterns.
-This often necessitates task-specific or data-driven heuristics, which can severely constrain their efficient generalization across diverse model architectures and datasets.
-Token merge, as a variant of sparse attention, merges the KV cache of similar tokens during inference to reduce memory usage zhang2024cam ; wang2024modeltellsmergeadaptive .
-Retrieval-based methods liu2024retrievalattention ; chenmagicpig ; zhang2024pqcache compress KV caches by retrieving relevant historical context for each query and are effective in tasks like RAG or document retrieval.
-However, they often lose substantial KV information, making them less suitable for general-purpose LLM inference.
-Model Compression Many model compression techniques closely align with KV cache compression in both goals and underlying methodologies.
+KV cache compression Beyond quantization, the field of LLMs is actively exploring advanced methods for KV cache compression.
+Sparse attention aims to reduce memory footprint by selectively handling the KV cache in a token-wise manner xiaoefficient ; chen2024prefixquant ; zhu2025oraclekv ; liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; li2024snapkvllmknowslooking .
+However, it discards the KV cache of a subset of tokens, even though the corresponding tokens may be required in subsequent decoding.
+Token Merging reduces memory usage by consolidating the KV caches of similar tokens during inference, achieving an effect related to sparse attention but through merging rather than dropping tokens zhang2024cam ; wang2024modeltellsmergeadaptive .
+Retrieval-based methods liu2024retrievalattention ; chenmagicpig ; zhang2024pqcache offload and index KV caches, retrieving a subset of relevant entries for each query, but introduce additional communication overhead.
+Model Compression Numerous model compression techniques share common objectives and methodological foundations with KV cache compression.
 GPTQ frantar2023gptqaccurateposttrainingquantization utilizes calibration set to reduce quantization induced degradation, while SmoothQuant xiao2024smoothquantaccurateefficientposttraining and AWQ lin2024awqactivationawareweightquantization minimize output error from the perspective of error propagation analysis.
-VQ-based methods such as VPTQ liu2024vptqextremelowbitvector and QUIP# tseng2024quipbetterllmquantization further enhance compression fidelity through residual VQ and Hadamard transform.
-Additionally, system-level frameworks like Atom zhao2024atomlowbitquantizationefficient and QServe lin2025qservew4a8kv4quantizationcodesign jointly optimize model and KV cache quantization, achieving end-to-end acceleration under low-bit quantization.
-Other works, such as Pruner-Zero dong2024prunerzeroevolvingsymbolicpruning and Parzc dong2024parzcparametriczerocostproxies , explore how to sparsify model weights while preserving model performance.
-Systems like FlashLLM xia2023flashllmenablingcosteffectivehighlyefficient and Spinfer fan2025spinfer leverage the model sparsity to accelerate inference.
+VQ-based methods such as QUIP# tseng2024quipbetterllmquantization further enhance compression fidelity through Hadamard transform.
+Pruner-Zero dong2024prunerzeroevolvingsymbolicpruning and Parzc dong2024parzcparametriczerocostproxies , explore how to sparsify model weights while preserving model performance.
+System-level works like Atom zhao2024atomlowbitquantizationefficient and QServe lin2025qservew4a8kv4quantizationcodesign Recent efforts jointly quantize model, KV cache and activatioin, enabling inference under low-bit and leveraging low-precision Tensor Cores to improve system performance, while approaches such as FlashLLM xia2023flashllmenablingcosteffectivehighlyefficient and Spinfer fan2025spinfer accelerate inference by leveraging model sparsity.
 3 Methodology
-Previous works on quantifying the KV cache typically do not distinguish the importance of individual tokens, but assume that all tokens contribute equally to attention accuracy, or they adopt a simpler strategy of only preserving designated sink tokens.
-However, our 1-bit VQ experiment on Mistral-7B contradicts this, unveiling a previously unaddressed aspect in KV cache quantization: a small subset of tokens provokes disproportionately large perturbations in the self-attention outputs.
-This divergence reveals a new opportunity for KV cache quantization.
-Instead of uniform application, the precision of quantization should correspond to each token’s individual importance.
-Specifically, if we identify and preserve a subset of high-impact tokens in FP16, we can unlock a far better memory–accuracy trade-off.
-Inspired by this, we present a novel token-aware VQ method for KV cache compression, namely, AnTKV.
-In AnTKV, we adopt a two-stage design: offline token-aware centroids learning (see Figure 3(a)) and online anchor token selection (see Figure 3(b)), the key of which lies in how to measure the importance of tokens.
-In the offline stage, we perform a weighted k-means clustering to produce centroids in VQ, and the gradient corresponding to each token is used as weights.
-In the online stage, to reduce the output error caused by quantization, we select tokens with both higher $\text{Err}_{\text{quant}}$ and AnS, i.e.
-anchor tokens, and keep them at FP16.
-Additionally, since applying RoPE disrupts the consistency of channel-wise magnitudes in $\bm{K}$ (see Appendix A for an illustrative example), which exhibits large inter-cluster distances and small intra-cluster variances, leading to large quantization error, the pre-RoPE strategy with the same as hooper2024kvquant is used in AnTKV.
-Figure 3: The overview of AnTKV.
-(a) In the offline stage, we perform token-aware centroid learning to obtain the centroids.
-(b) In the online stage, we apply VQ using centroids and compute the AnS.
-Then identify the top-$k$ anchor tokens with the highest product of VQ error and AnS, and retain them in full-precision to preserve model performance.
+VQ is employed as it enables sub-bit compression, which SQ cannot achieve, and in the ultra-low-bit regime ($\leq 2$ bit) it leverages intra-vector correlations to attain significantly better quantization quality compared to SQ.
+As shown in the figure 1, we observe that although per-token KV cache quantization errors appear similar, the resulting attention output errors differ dramatically.
+Moreover, the error distribution is steep, with large errors occurring unpredictably across tokens.
+Previous works on KV cache quantization typically treat all tokens equally or are biased toward attention sinks, as inspired by AttentionSink xiaoefficient , which we argue is suboptimal.
+Our observation reveals a new opportunity for KV cache quantization, where tokens with greater influence on attention output error should be prioritized.
+Building on this observation, we propose AnTKV, a token-aware VQ framework for KV cache quantization.
+It adopdts a dual-stage design consisting of offline token-aware centroid learning (Figure 3(a)) and online anchor token selection (Figure 3(b)).
+The central challenge lies in effectively measuring token importance.
+During centroids leanring, forward error analysis provides per-token error propagation factors, which serve as weights in a token-aware $k$-means clustering to generate VQ centroids, thereby emphasizing tokens with significant influence on attention-output error.
+For anchor token selection, directly using error propagation factors incurs expensive gradient computation.
+To avoid this cost, we introduce AnS, a lightweight metric derived from attention forward error analysis that enables efficient prompt-aware identification of anchor tokens.
+Figure 3: Overview of AnTKV.
+In the stage (a), token-aware centroids are learned from calibration data through weighted clustering, where the weights are error-propagation factors obtained by forward error analysis.
+In the stage (b), the KV cache is quantized with centroids, and AnS is computed to identify anchor tokens, which are preserved in full precision to mitigate accuracy loss.
 3.1 Offline Token-Aware Centroids Learning
-The offline stage aims to learn centroids by aggregating KV cache from general calibrations.
+In this stage, centroids are learned offline by aggregating KV cache from calibration data.
 Specifically, each row (per token) of $\bm{K}$ and $\bm{V}$ is divided into several sub-vectors, then using $k$-means to cluster them, and only the centroids are stored.
-To integrate the importance of tokens, we introduce the weighted k-means instead of classical k-means for the offline centroids learning.
-From the perspective of accuracy, we hope that the errors caused by quantizing $\bm{K}$ and $\bm{V}$ will not affect the accuracy of LLMs as much as possible. Mathematically, it can be formulated as
-$\begin{split}\min\limits_{\bm{C}^{k}\in\mathbb{R}^{c^{k}\times d}}|L(\bm{K},\bm{V})-L(\texttt{vq}(\bm{K}),\texttt{vq}(\bm{V}))|,\\ \end{split}$ (1)
+Specifically, each token row of $\bm{K}$ and $\bm{V}$ is partitioned into sub-vectors and clustered with $k$-mean.
+The centroids are retained for online KV cache quantization.
+From an accuracy perspective, replacing $\bm{K}$ and $\bm{V}$ with their corresponding centroids should introduce minimal impact on model output. Mathematically, it can be formulated as
+[…]
 where $c^{k}$ is the number of clusters, and $\bm{C}^{k}$ is the matrix consisting of centroids. By the first-order Taylor series expansion, we have
@@
 where $\frac{\partial L}{\partial\bm{K}_{j,:}}$ and $\frac{\partial L}{\partial\bm{V}_{j,:}}$ are the gradient of the loss function $L(\cdot)$ with respect to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$, i.e., the $j$th row of $\bm{K}$ and $\bm{V}$ that correspond to the $j$th token.
-From (1) and (2), the weighted k-means clustering with gradients as weights is essentially to find a clustering strategy that minimizes the error caused by KV quantization.
-This also explains the effectiveness of the weighted k-means clustering used in this stage.
+From equation (1) and (2), the weighted k-means clustering with gradients as weights is essentially to find a clustering strategy that minimizes the error caused by KV quantization.
 3.2 Online Anchor Token Selection
-In the online stage, the gradient can no longer be used as AnS due to prohibitive memory and computational costs under real-time constraints.
-To this end, we here give the error propagation analysis of the self-attention (i.e., Attn).
-Specifically, the error propagation analysis is based on the perturbation bound of Attn that is related to each row of $\bm{K}$ and $\bm{V}$; see Theorem 3.1 (see Appendix B for proof).
+In online inference, gradients can no longer serve as the token importance metric due to the cost imposed by real-time constraints.
+To address this, we perform an error propagation analysis of the attention operator (Attn).
+Specifically, the analysis derives a perturbation bound of Attn with respect to each row of $\bm{K}$ and $\bm{V}$, as presented in Theorem 3.1, with the detailed proof provided in Appendix B.
 Theorem 3.1.
@@
 where $\bm{e}\in\mathbb{R}^{n}$ is a vector whose entries are all $1$.
-We remark that the error propagation factors corresponding to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$ given in Theorem 3.1 can be regarded as the upper bound of the gradient of the self-attention operator related to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$.
-The computation involving $\bm{K}$ in Theorem 3.1 entails substantial overhead, rendering it impractical for use during online inference.
-To address this, we propose a simplified variant that excludes the contribution of $\bm{V}$ to the quantization error of $\bm{K}$.
-By this simplification, we obtain the following reformulation.
-A detailed evaluation of the effectiveness of AnS is provided in Appendix C.
-$\text{AnS}(\bm{K}_{j,:})=\sum_{i}\bm{A}_{i,j}(1-\bm{A}_{i,j})\cdot\left\|\bm{Q}_{i,:}\right\|_{2}$ (5)
-During the prefill phase, AnS serves as an effective metric to identify anchor tokens that preserve critical contextual information by measuring their impact on attention outputs.
-In contrast, although AnS can still be computed during decoding, the anchor token it identifies may have already been quantized during prefill, preventing the preservation of its FP16 value for reducing accumulated errors.
-An important observation is that $\bm{A}$ and $\bm{A}\odot(\bm{E}-\bm{A})$ exhibit strong locality during decoding (see Appendix D), with anchor tokens primarily concentrated at the beginning and end of the sequence.
-In particular, the beginning tokens—referred to as sink tokens—are almost always identified as anchor tokens during prefill, indicating that our method effectively handle sink tokens.
-Motivated by this locality, we adopt a sliding window approximation of AnS during decoding.
+We remark that the error propagation factors corresponding to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$ given in Theorem 3.1 can be regarded as the upper bound of the gradient of the attention operator related to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$.
+The computation involving $\bm{K}$ in Theorem 3.1 introduces significant overhead and is therefore unsuitable for online inference.
+To address this limitation, we propose a simplified variant that excludes the contribution of $\bm{V}$ to the quantization error of $\bm{K}$.
+This leads to the following reformulation, with a detailed evaluation of AnS effectiveness provided in Appendix C.
+$\text{AnS}(\bm{V}_{j,:})=\sum_{i}\bm{A}_{i,j}\quad\quad\text{AnS}(\bm{K}_{j,:})=\sum_{i}\bm{A}_{i,j}(1-\bm{A}_{i,j})\cdot\left\|\bm{Q}_{i,:}\right\|_{2}$ (5)
+In online inference, during the prefill phase, AnS serves as an effective metric for identifying anchor tokens that induce substantial accuracy loss.
+In autoregressive decoding phase, AnS can still be computed.
+However, the anchor tokens it identifies may already have been quantized, which prevents preserving their full-precision values and limits error reduction.
+An important observation is that both AnS(K) and AnS(V) exhibit strong locality during the decoding phase (see Appendix D), with anchor tokens predominantly concentrated at the head and tail of the sequence.
+Experimental results show that the anchor tokens at the head of the sequence are consistently identified during the prefill stage, corresponding to sink tokens.
+This observation further demonstrates the effectiveness of our method.
+Building on the tail locality, we use a sliding-window approximation of AnS during decoding to further enhance efficiency while mitigating accuracy degradation.
 3.3 Implementation
-For the offline stage, the gradient of $\bm{K}$ and $\bm{V}$ is used as a weight to learn centroids.
+For the offline centroids learning stage, the gradients of $\bm{K}$ and $\bm{V}$ are employed as weights for centroid learning.
 We implement it with a custom LinearWithAct to capture KV cache and corresponding gradients.
 Subsequently, we employ the weighted k-means provided by cuML to perform efficient clustering.
-For the online stage, AnS is derived from the error propagation factor given in Theorem 3.1 and (5).
-We design and implement a dedicated GPU kernel using Triton to efficiently compute it in conjunction with FlashAttention, enabling long-context inference.
-Since AnS requires reduction operations over the attention score matrix $\bm{A}$ and its transformed form $\bm{A}\odot(\bm{E}-\bm{A})$ along query (column-wise), directly fusing its computation into FlashAttention is impossible.
-To preserve the efficiency of FlashAttention, we decouple the AnS computation and execute it immediately afterward.
-To support this design, we extend FlashAttention to preserve three auxiliary tensors: the $L_{2}$ norm of each query vector, the key-wise (row-wise) sum, and the key-wise maximum of the matrix $\bm{Q}\bm{K}^{T}$.
-These additional outputs enable efficient reconstruction of the attention scores and allow AnS to be computed with minimal overhead.
-More details about the implementation are described in Algorithm 1 of Appendix E.
+In the online stage, AnS is derived from the error propagation factor given in Theorem 3.1 and Equation (5).
+To enable efficient long-context inference, we design and implement a dedicated GPU kernel using Triton that computes AnS in conjunction with FlashAttention.
+Because AnS requires reduction operations over the attention score matrix $\bm{A}$ and its transformed form $\bm{A}\odot(\bm{E}-\bm{A})$ along the query dimension (column-wise), direct fusion into FlashAttention is infeasible.
+To preserve the efficiency of FlashAttention, we decouple AnS computation and execute it immediately afterward.
+For this purpose, we extend FlashAttention to additionally output three auxiliary tensors: the $L_{2}$ norm of each query vector, the key-wise (row-wise) sum, and the key-wise maximum of the matrix $\bm{Q}\bm{K}^{T}$.
+These tensors allow the reconstruction of the attention scores and facilitate AnS computation with minimal overhead.
+Further implementation details are provided in Algorithm 1 of Appendix E.
+Finally, since the application of RoPE disrupts the channel-wise magnitude distribution of $\bm{K}$ (see Appendix A), which otherwise exhibits large inter-cluster distances and small intra-cluster variances, the pre-RoPE strategy, consistent with hooper2024kvquant , is adopted in AnTKV.
 4 Experiments
-In this section, we compare the performance of AnTKV with existing KV quantization methods through extensive experiments.
-The experimental setup, including models, benchmark datasets, evaluation metrics, and selected baselines, is as follows.
+In this section, we present an extensive comparison between AnTKVand existing KV quantization methods. The experimental setup is detailed below.
 Models, Datasets, Metrics, and Parameter Settings.
-To validate the effectiveness of AnTKV in KV cache quantization and ensure the generality of results, we evaluate five representative models from the LLaMA and Mistral families: LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-8B-Instruct, Mistral-7B, and Mistral-7B-Instruct-v0.3.
-We use 128 samples of length 2048 from the WikiText2-v1 training set as the calibration.
-Perplexity is evaluated on WikiText-2 and C4, while zero-shot accuracy on MMLU hendrycks2021mmlu , ARC-C clark2018arc , MathQA amini2019mathqa , and PIQA bisk2020piqa is used to assess understanding and reasoning.
-Long-context performance is measured using LongBench bai2024longbench .
-For perplexity and zero-shot benchmarks, quantized KV cache is employed for attention.
-In LongBench, the FP16 KV cache is first utilized to compute attention outputs, before quantization is applied.
-[…]
+To validate the effectiveness and generality of AnTKVin KV cache quantization, we evaluate five representative models from the LLaMA and Mistral families.
+For calibration, 128 samples of length 2048 are drawn from the WikiText2 training set.
+Model quality is assessed through three categories of benchmarks: (i) perplexity on WikiText-2 and C4; (ii) zero-shot accuracy on MMLU hendrycks2021mmlu , ARC-C clark2018arc , MathQA amini2019mathqa , and PIQA bisk2020piqa to evaluate understanding and reasoning; and (iii) long-context performance on LongBench bai2024longbench .
+For perplexity and zero-shot evaluations, quantized KV caches are directly used for attention outputs, whereas for LongBench, full precision KV cache is used to compute attention outputs and quantized KV cache is used during decoding.
+Across all benchmarks, anchor tokens are restricted to a small subset: 1% of the context length for perplexity, 16 for understanding and reasoning, and 64 for LongBench.
+For fair comparison, a sliding window of size 32 is applied in LongBench, following the mainstream setting.
+Baselines.
+We compare AnTKVwith full precision and representative KV cache quantization methods, including KIVI liu2024kivi , SKVQ duanmu2024skvq , KVQuant-1% hooper2024kvquant , and CQ zhang2024kv .
+SKVQ is configured with a group size of 64 and five sink tokens xiaoefficient , while KVQuant retains four sink tokens.
+Since CQ results are not publicly available, we reproduced them following the methodology in their paper to the best of our understanding.
+For VQ settings, we adopt the notation “dnncmm”, covering 4-bit (d2m256), 2-bit (d4m256), 1-bit (d8m256), 0.75-bit (d16m4096), and 0.375-bit (d32m4096).
 4.1 Perplexity Results
-Perplexity is a standard metric that is widely used to evaluate the quality of the output of LLMs, with lower values indicating better performance.
+Perplexity is a standard benchmark that is widely used to evaluate the quality of the output of LLMs, with lower values indicating better performance.
 The perplexity results for different KV quantization approaches on WikiText-2 and C4 are presented in Table 1.
 The results in this table indicate that the proposed AnTKVconsistently achieves competitive or superior perplexity across various bit-widths and model architectures.
-Under 4-bit and 2-bit quantization, AnTKV achieves competitive or superior performance compared to existing approaches.
-In the 1-bit and sub-bit regimes, AnTKV significantly outperforms all baselines.
-It is worth mentioning that on the C4 dataset under sub-bit quantization, the baseline methods exhibit extremely high perplexity due to anchor tokens not being well captured by fixed centroids.
-Thanks to effective AnS design and innovative Anchor Token selection, AnTKV substantially reduces perplexity, e.g., from 66.28 to 14.42 (LLaMA-3-8B, 0.75-bit).
-Table 1: The evaluations are conducted using each model’s maximum context length, specifically 4096 for LLaMA-2-7B and 8192 for LLaMA-3-8B and Mistral-7B.
-"Ours" refers to our method (AnTKV) without using any anchor tokens, while Ours-1% denotes AnTKV where 1% of the tokens are treated as anchor tokens and retained in FP16.
-For simplicity, the reported bit-widths do not consider the effect of centroids.
+Under 4-bit and 2-bit quantization, it achieves competitive performance compared to baseline.
+In the 1-bit and sub-bit regimes, it significantly outperforms all baselines.
+On the C4 dataset under sub-bit quantization, baseline methods suffer from extremely high perplexity, as fixed centroids fail to capture anchor tokens.
+By contrast, with its effective AnS design and anchor token selection, AnTKVsubstantially lowers perplexity, reducing it from 66.28 to 14.42 on LLaMA-3-8B at 0.75-bit.
+Table 1: All evaluations are performed under the maximum context length of each model, specifically 4096 for LLaMA-2-7B and 8192 for LLaMA-3-8B and Mistral-7B.
+“Ours” refers to AnTKVwithout anchor tokens, whereas “Ours-1%” denotes AnTKVwith 1% of tokens designated as anchor tokens and retained in FP16.
+For clarity, the reported bit-widths exclude the contribution of centroids.
 Bit LLaMA-2-7B LLaMA-3-8B Mistral-7B
@@
 4.2 Understanding and Reasoning Benchmark
-To ensure that AnTKV exhibits broad understanding and reasoning capabilities, we evaluate its performance on four representative benchmarks: MMLU, ARC-Challenge, PIQA, and MathQA, using both LLaMA-8B-Instruct and Mistral-7B-Instruct.
-These benchmarks respectively assess multi-domain knowledge reasoning (MMLU), complex question answering (arc_challenge), commonsense reasoning (PIQA), and mathematical problem-solving (MathQA).
-Figure 4 illustrates the comparison results of AnTKV with SKVQ and CQ.
-Similarly, it can be seen that AnTKV consistently outperforms SKVQ and CQ across all four benchmarks under low-bit quantization, especially at 1-bit and sub-bit levels.
-Figure 4: Understanding and Reasoning ability evaluation accuracy results on MMLU, ARC-C, PIQA, and MathQA under different quantization bit-widths.
+To assess the breadth of AnTKV’s understanding and reasoning capabilities, we evaluate it on four representative benchmarks using LLaMA-8B-Instruct and Mistral-7B-Instruct.
+These benchmarks target multi-domain knowledge reasoning (MMLU), complex question answering (ARC-Challenge), commonsense reasoning (PIQA), and mathematical problem solving (MathQA).
+Due to missing the implementations in the official repository, KIVI and KVQuant are not included.
+As shown, AnTKVconsistently maintains higher accuracy across LLaMA and Mistral models, with particularly strong advantages at 1-bit and sub-bit settings where baseline methods degrade sharply.
+Figure 4: Evaluation of understanding and reasoning accuracy on MMLU, ARC-C, PIQA, and MathQA under different quantization bit-widths.
 4.3 Long-Context Benchmark
@@
 Figure 5 shows that AnTKV consistently outperforms CQ.
-Notably, when quantizing down to 0.375-bit, AnTKV maintains a tolerable performance degradation, with the average score dropping from 46.2 to 34.5.
+Notably, when quantizing down to 0.375-bit, AnTKV maintains a tolerable performance degradation, with the average score dropping from 46.2 to 38.1.
 Figure 5: The evaluation accuracy results on LongBench under different KV cache quantization bit-widths.
 4.4 Efficiency
-In this experiment, we evaluate the efficiency of our AnTKV implementation against the FlashAttention-accelerated FP16 baseline on LLaMA-3-8B using a single A100-80GB GPU.
+In this experiment, we evaluate the efficiency of our AnTKV implementation compared with huggingface baseline transformers on LLaMA-3-8B using a single A100-80GB GPU.
 As shown in Figure 6,  AnTKV substantially extends the maximum context length from 128K to 384K.
-In long-context scenarios, we observe that intermediate activations consume a significant portion of memory.
-By applying a series of in-place operations, AnTKV ultimately supports up to 810K tokens under 1-bit quantization and 840K under 0.375-bit quantization.
-More importantly, AnTKV sustains low memory consumption while maintaining acceptable latency.
-Under 1-bit quantization, a noticeable increase in prefill latency is observed only at the 128K context length, which represents the upper limit supported by most current models.
-It is observed that 0.375-bit quantization introduces higher latency than its 1-bit.
-This increase in latency arises from the use of a larger number of centroids, which are employed to enhance model accuracy.
-To evaluate decoding efficiency, we evaluate the throughput of AnTKV under a fixed context length of 1K tokens.
-As shown in Figure 7, AnTKV supports significantly larger batch sizes compared to the baseline method.
-By substantially reducing KV cache access, it improves throughput across all bit-widths.
-Particularly, the maximum throughput reaches 3.5x that of the baseline method under 1-bit quantization.
-Figure 6: Memory and latency comparison. Bars show memory usage (red: 1-bit, green: 0.375-bit); lines indicate prefill latency across context lengths.
-Figure 7: Decoding throughput comparison. Our method supports larger batch sizes and achieves higher throughput.
+In long-context inference, our profiling shows that intermediate activations account for a substantial portion of memory usage.
+By introducing a series of in-place operators, AnTKV supports up to 810K tokens under 1-bit quantization and 840K under 0.375-bit quantization, while maintaining low memory consumption.
+To evaluate decoding efficiency, we measure the throughput of AnTKV with a fixed context length of 1K tokens.
+As shown in Figure 7AnTKV enables substantially larger batch sizes and improves throughput across all bit-widths by reducing KV cache access.
+In particular, under 1-bit quantization, the maximum throughput reaches $3.5\times$.
+Figure 6: KV cache memory size comparison. Gray bars denote full precision, red bars 1-bit, and green bars 0.375-bit quantization. Striped bars indicate results with in-place operators enabled.
+Figure 7: Decoding throughput comparison. Our method supports larger batch sizes and achieves higher throughput. Dashed lines indicate results with in-place operators enabled.
 4.5 Ablation Study
@@
 5 Limitation & Conclusion
-Although AnTKV demonstrates its advantages in experiments, it also has several limitations.
+Although AnTKV demonstrates its advantages in experiments, it also has few limitations.
 First, more accurate AnS for tokens and higher performance implementations for its computation may be possible.
@@
 Nevertheless, further empirical validation is required.
-We observed from several experiments that different tokens exhibit varying sensitivity to KV quantization, which opens a new opportunity for a better memory-accuracy trade-off.
-To this end, we present AnTKV, a two-stage token-aware VQ framework that involves offline gradient weighted centroid learning and online anchor token selection guided by AnS.
-Moreover, to ensure the efficiency of AnTKV in LLM inference, we implement a lightweight GPU kernel compatible with FlashAttention to compute AnS with minimal overhead.
-Extensive experiments show that AnTKV achieves superior performance under low and ultra-low-bit quantization, significantly extending the maximum context length (up to 840K tokens) on a single A100 GPU while maintaining low perplexity and high throughput.
-Our results demonstrate the potential of token-aware quantization in scaling LLM inference to longer contexts without sacrificing accuracy.
+This work addresses the preservation of accuracy under ultra low bit KV cache quantization.
+We propose AnTKV , a vector quantization based framework that exploits intra vector correlations.
+AnTKV uses a dual stage design with offline token aware centroid learning and online anchor token selection, which mitigates the disproportionate error from anchor tokens.
+Across the LLaMA and Mistral families, AnTKV attains accuracy close to full precision and consistently surpasses baselines in the ultra low bit regime.
+It also scales LLaMA-3-8B to 840K tokens on a single 80 GB A100, and increases decoding throughput by up to $3.5\times$.
 Appendix A Distribution of pre- and post- RoPE Key
@@
 Further, we can obtain
-$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K}+\delta\bm{K},\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}&\approx\|\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Diag}\left(\texttt{Vec}(\bm{A})\right)\texttt{Vec}(\bm{Y})\|_{L_{1}}\\ &=\sum\limits_{i}\|\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\bm{X}_{i,:}^{T}\|_{L_{1}}\\ &\leq\sum\limits_{i}\sum\limits_{j}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}|\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}|\\ &\leq\sum\limits_{j}\sum\limits_{i}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}\|{\bm{Q}}_{i,:}\|_{2}\|{\delta\bm{K}}_{j,:}\|_{L_{1}}.\\ \end{split}$ (9)
+[…]
 For $\bm{V}$, we have
@@
 2:Output: AnS of KV, i.e., $\text{AnS}({\bm{K}})$ and $\text{AnS}({\bm{V}})$
-3:  $(\bm{O},\bm{L},\bm{M},\left\|\bm{Q}_{i,:}\right\|_{L_{2}})\leftarrow\texttt{FlashAttention}(\bm{Q},\bm{K},\bm{V})$
+[…]
 4:for each block key index $j$ in parallel (assigned to GPU block) do
@@
 7:   $\bm{A}_{i,j}\leftarrow\exp(\bm{S}_{i,j}-\bm{M}_{h,i})/\bm{L}_{h,i}$
-[…]
-[…]
+8:   $\text{AnS}({\bm{K}})_{i}\leftarrow\text{AnS}({\bm{K}})_{i}+\texttt{col\_sum}(\bm{A}_{i,j}\cdot(1-\bm{A}_{i,j}),\ \text{row-wise})$
+9:   $\text{AnS}({\bm{V}})_{i}\leftarrow\text{AnS}({\bm{V}})_{i}+\texttt{col\_sum}(\bm{A}_{i,j},\ \text{row-wise})$
 10:  end for
@@
 The results are shown in Figure 13.
-The results indicate that Trec’s performance increases linearly with the number of anchor tokens, eventually reaching a result within 10 points of the baseline when using $10\%$ anchor tokens.
-In contrast to Trec, Qasper’s accuracy remains stable with a small number of anchor tokens but shows a significant performance improvement when increased to $10\%$.
-We believe this difference stems from the inherent properties of each task.
+it illustrates that both the Trec and Qasper subtasks exhibit a consistent improvement pattern as the number of anchor tokens increases.
+In particular, moving from 0% to 1% anchor tokens leads to a substantial performance gain across all quantization settings, highlighting that even a very small proportion of anchor tokens can effectively mitigate the degradation introduced by ultra-low-bit quantization.
+Beyond this point, the improvements from 1% to 2%, 2% to 5%, and 5% to 10% follow an approximately linear trend, with performance gradually approaching the FP16 baseline.
+These results demonstrate that anchor tokens play a dual role that a small fraction is sufficient to deliver immediate and significant benefits, while larger allocations further provide steady, near-linear enhancements in downstream task performance.
 Figure 13: Trec and Qasper results on LLaMA-3-8B-Instruct with varying anchor token numbers.